OpenStudy (pottersheep):

Grade 11 chem help please~ What mass of amoonia gas will be deco,posed when 10.0 moles of hydrogen are formed at 300 degrees CELCIUS and 795 mmg pressure? What does the whole ideal gas law have to do with the mass?? Btw, the answer is supposed to be 1.13 x 10^2. Thank you!!

sam (.sam.):

Actually, the temperature, pressure has nothing to do with it unless its an equilibrium reaction. We'll setup the reaction first. \[2NH_3 \leftarrow \rightarrow N_2 +3H_2\] We have 10.0 moles of H2, using the fraction method, \[10.0~\text{moles of }H_2 \times \frac{2~mol NH_3}{3~mol H_2} \frac{17~gNH_3}{1~molNH_3}\] \[10.0~\cancel{\text{moles of }H_2} \times \frac{2~\cancel{mol NH_3}}{3~ \cancel{mol H_2}} \frac{17~gNH_3}{1~\cancel{molNH_3}}=113 \text{ g of }NH_3=1.13 \times 10^2~g\]

OpenStudy (pottersheep):

ohhhhhhhhh

OpenStudy (pottersheep):

Thank you!!!!